7) a) If the corporate tax rate is 40% and the individual tax rate is 20% (and these are the only taxes), if a corporation has $50,000 of operating income, what is the after tax income for its owners? The tax rate is 15% between $0-$50,000 and 25% between $50,001-$100,000. If you earn $60,000, then what is your average and marginal tax rates? A bank has quoted to you a monthly rate of 0.25% on a short-term loan. What is the stated annual rate of that loan? c) d) What is the future value of $8 in three years if the interest rate is 20% and the compounding is continuous? e What is the future value of a 3-year annuity due with $50 cash flows if the annually compounded interest rate is 6%? f)If we need to find the future value of a current cash flow and interest is compounded quarterly, how many periods of compounding occur in two and half years? g What is the difference between the present value of a 3-year ordinary annuity with a S20 cash flow and the present value of a 3-year annuity due with a S20 cash flow when the interest rate is 3.5%?
Answer #1

As per rules I am answering the first 4 subparts of the question

A: After tax income for the owners = Operating income*(1-corporate tax)*(1- personal tax)

= 50000*(1-0.4)*(1-0.2)

=24000

B: Marginal tax rate is the highest 25%

Total tax payable = (60000-50000)*25% + 50000*15% = 10000

Average tax = tax/ Income

=10000/60000

= 16.67%

C: Annual rate = monthly rate * 12

= 0.25%*12

= 3%

D: Fv = PV*e^rt

= 8* 2.718^(0.2*3)

=$14.57